# Displacement Memory Effect from Supersymmetry

### v1

@@ 3.2 DME of plane GWs
Now, it is time for our second goal. Let us begin with $n=1$. Either from (A.7) or (A.8), we get
@@
Provided that $A\geq 1$, the velocity of $X_{1}(U)$ is vanishing at infinities
$\frac{dX_{1}}{dU}|_{U\to\pm\infty}=0,$ (3.8)
so that it exhibits DME [^2] It is a half wave in the language of [9, 10, 11]. One can also see that at distant times $U=\pm\infty$, $X_{1}(U)\neq 0$ only if $A=n=1$.
Subsequent SUSY Darboux transformations (A.7) generate other geodesics with higher ”half-wave numbers” $n$.
With this unified approach, we will show below that the PT and Scarf profiles in [9, 10, 11] are actually two different faces of the same theory from the perspective of SUSY QM. Both profiles are rooted in (3.1), hence derivable from (3.2).
One important point is the value of the free parameter $A\geq n$. If we set $A=n$, the formula for geodesics (A.8) simplifies in terms of $W(U)$ as
@@
If we substitute (3.9) into (3.10), we obtain
$\Delta X_{n}$ $=$ $X_{n}(\infty;n)-X_{n}(-\infty;n)$ (3.11)
$=$ $W(\infty;n)X_{n-1}(\infty;n-1)-W(-\infty;n)X_{n-1}(-\infty;n-1),$ (3.12)
as our geodesics satisfy the Neumann boundary condition (2.5). Therefore, SUSY QM yields a nice formula (3.12) for the DME of plane waves in terms of asymptotic values of superpotentials and preceding geodesics333(3.12) might be somewhat reminiscent of Christodoulou’s nonlinear memory [25].. While (3.11) shows that $\Delta X\neq 0$ condition excludes parity even geodesics, second formula (3.12) indicates that DME is related to the parity of superpotential and the preceeding geodesics.
But individual parts of our superpotential (3.1) have different parity properties. Thus, we will investigate each part separately and show that these magical numbers come out of our unified picture in a beautiful manner.
3.2.1 Pöschl-Teller profile
To confine ourselves with the PT type profile (2.6), we need to set
@@ 3.2.1 Pöschl-Teller profile
$A-n$ $=0$ (3.13b)
in (3.5). While the first condition eliminates the parity-odd scarf term, the second one selects the highest excited state by quantizing the parameter $A$. Selecting $A$ as an integer leads to the so-called reflectionless PT problem. In that case, the geodesic equation turns out to be
@@
cf.(2.2a) As we are dealing with a shape invariant profile, the bounded geodesics can be found via (A.7), where we set $A=m$ first,
@@
where $X_{0}(U,m)=\operatorname{sech}^{m}U$ and $m=1,2,3,..$.
@@
For instance, if we take $n=1$, we obtain
@@
$X_{1}=\tanh U$ (3.17)
in line with [9, 10]. DME effect can be computed very easily from (3.12) as
@@
Similarly, we can obtain $n=2$ geodesic using (3.15) as
@@
For $n=3$, solution of the geodesic equation (3.14) is
$X_{3}(U;m)=(2m-5)(2m-3)\Big{(}2(m-2)\operatorname{sech}^{m-3}U+(1-2m)\operatorname{sech}^{m-1}U\Big{)}\tanh U$ (3.21)
@@
One can go on and verify that $\Delta X_{n}(m)\neq 0$ only for $m=n$ and $n$ is an odd integer. We conclude that the fine-tuned parameters yielding DME are nothing but the highest excited states of the analogous quantum mechanical problem satisfying the Neumann boundary conditions.
Longitudinal motion and memory
In order to compute $V$-component of geodesics for the first two $n$, we substitute our solutions for $X(U)$ i.e., (3.17) and (3.20) into (2.3) respectively. We observe that part of the longitudinal motion organizes itself as another particular solution of the PT problem. For instance,
@@
$V_{2}(U)$ $=$ $-3\tanh U\operatorname{sech}^{2}(2-3\operatorname{sech}^{2}U)+c_{3}U+c_{4}=-\frac{1}{35}X_{3}(U;m=5)+c_{3}U+c_{4},\qquad$ (3.24)
where $c_{1},..,c_{4}$ are constants. $X_{1}(U;m=3)$ and $X_{3}(U;m=5)$ can be found from (3.16) and (3.21), respectively. As $n\neq m$, those motions exhibit VME provided that initial constants $c_{1},c_{3}$ do not vanish.
3.2.2 Scarf profile
In order to reproduce the geodesic equation in [11] for scarf profile (2.7), we need to satisfy another set of equations
@@ 3.2.2 Scarf profile
in (3.5). While the first equation quantizes the parameter $A$, the second one solves for $B$ as
@@
where $n=0,1,2,\cdots$. Pay attention to the fact that $B$ has double roots. This multiplicity will allow us to extend our arguments to exact plane gravitational waves with 2-transverse directions [^4] That extension was first done in [11]. Here, we elaborate on their results using SUSY methods. in Section 4.
In $1$-transverse dimensional case, (3.5) reduces to a geodesic equation (2.2)
@@
augmented with (2.5). Thus, supersymmetry leads to a very quick derivation of the scarf geodesic equation that mimics the derived Gaussian flyby profile. Note that, in [11] above equation was obtained through the hypergeometric equation and solved with the Nikiforov-Uvarov method. SUSY provides a more economical way for the solutions.
@@
The seed geodesic can be found via (A.4) as
$X_{0}=\operatorname{sech}^{m}U\ e^{-B\tan^{-1}\sinh U},$ (3.28)
where $A=m$ and $B=\pm\sqrt{n(n+1)}$. As before, $m$ is a positive integer. When $n=1$ geodesic can be obtained from (A.7)
@@
It is only when $m=n=1$ (3.29) yields a non-vanishing asymptotic value. In this case, it reduces to
@@
Let us emphasize that (3.30) satisfies (3.27) with $n=1$. The related DME can be computed either directly from the geodesic solution or from (3.12)
$\Delta X_{1}=2\cosh\left(\frac{B\pi}{2}\right)=2\cosh\left(\frac{\pi}{\sqrt{2}}\right),\quad B=\pm\sqrt{2}$ (3.31)
@@
Applying a second Darboux transformation, we generate $X_{2}(U;m)$. In order to satisfy (3.27), we immediately set $m=2$ and get
@@
By virtue of supersymmetric transformations (3.9), one may obtain geodesics with non-zero DME in a straightforward manner.
Longitudinal motion and memory
In order to illustrate the longitudinal memory, we may construct $V_{1}(U)$ (2.3)
@@
As
$\frac{dV_{1}}{dU}|_{\pm\infty}=c_{1},$ (3.36)
where the coordinate $V$ exhibits a VME. This result is generic for any $V_{n}(U)$.
@@
Apart from those 3-dimensional cases presented above, we may construct plane waves with 2-transverse directions [^5] One may try combinations of Scarf and PT profiles too.. For this, we choose a diagonal profile $K$ in (2.1), set $X^{i}=\{X,Y\}$ and begin with Scarf-type profile first.
4.1 Exact plane gravitational wave with Scarf profile
The double root of $B$ (3.26) allows Zhang et al. [11] to construct an exact plane gravitational wave (vacuum solution of Einstein’s equations) mimicking the derivative Gaussian profile (2.4). The associated profile is diagonal
$K_{scarf}=\begin{pmatrix}(2n+1)|B|\operatorname{sech}U\tanh U&0\\ 0&-(2n+1)|B|\operatorname{sech}U\tanh U\,\end{pmatrix}$ (4.1)
and traceless. $|B|=\sqrt{n(n+1)}$ and $n=1,2,3,...$.
First, 2 sets of geodesics with their DMEs can be simply extracted from Section 3.2.2 as
@@ 4.1 Exact plane gravitational wave with Scarf profile
$Y_{1}(U)$ $=\left[\tanh U-2\sqrt{2}\operatorname{sech}U\right]e^{\sqrt{2}\tan^{-1}(\sinh U)},\quad\quad\Delta Y_{1}=2\cosh\left(\frac{\pi}{\sqrt{2}}\right),$ (4.2b)
@@
$Y_{2}$ $=\left[2+21\operatorname{sech}^{2}U-8\sqrt{6}\tanh U\operatorname{sech}U\right]e^{\sqrt{6}\tan^{-1}(\sinh U)},\quad\Delta Y_{2}=4\sinh\Big{(}\frac{\sqrt{6}\pi}{2}\Big{)}.$ (4.3b)
The above results are in line with the ones in [11].
It is also possible to find lightcone geodesic $V_{n}(U)$ by combining solutions of transverse geodesics. As explained in Section 3.2.2, the longitudinal memory will be a VME.
We may construct further plane wave solutions where $m>n$, see (3.29). In that case, the trace of the profile is non-zero, i.e., $\mathrm{Tr}{K}=2(m-n)^{2}$. One can get a non-vanishing DME when, for instance, $n=2$ and $m=3$.
@@ 4.2 Pöschl-Teller plane waves
In [10], it was argued that when an exact plane GW is built out of a PT profile, DME can be obtained only in one direction. The other direction should be put to zero by initial conditions. SUSY provides a reasoning for this: Unlike the scarf case, the parameter $A$ of the PT profile has a single root (3.13b).
To understand this better, we may construct a 4-dimensional plane wave out of supersymmetric partners (3.2). The result will be a null-fluid spacetime, $\mathrm{Tr}(K)\neq 0$, with the following diagonal profile
$K_{PT}=\begin{pmatrix}(m-n)^{2}-m(m+1)\operatorname{sech}^{2}U&0\\ 0&(m-n)^{2}-m(m-1)\operatorname{sech}^{2}U\,\end{pmatrix}$ (4.4)
where parameters $m$ and $n$ are integers. From Section (3.2.1), we recall that DME occurs only when $m=n$. That simplifies (4.4) and puts it into its final form. Then, resulting geodesic equations (2.2a) are
@@
$\left(\frac{d^{2}}{dU^{2}}+m(m-1)\operatorname{sech}^{2}U\right)Y$ $=0$ (4.5b)
Due to shape invariance, the solutions of $X$ and $Y$ are interrelated: they match with each other except the ground state of $X$. If one sets $m=1$, the solution for the first one is $X=\tanh U$ (3.17). The second equation, on the other hand, becomes that of a free particle and confirms the claim made in [10]. Recall that choosing $m$ as an integer leads to a ”reflectionless PT potential” whose reflection coefficient is vanishing just like a free particle.
Other choices of $m>1$ yield DME again in one direction. For instance, when $m=2$ $X=3\tanh^{2}U-1$ and $Y=\tanh U$. Therefore, the usage of superpartners in the profile yields a tower of geodesics exhibiting DME in alternating directions.
@@ 5 Modeling singular profiles with SUSY
It is time for our third goal. We show that SUSY provides further insight and predicts DME for other GWs.
In [26, 27], memory effects of the profile $K=\frac{1}{U^{4}}$ was discussed. Later, authors in [7] obtained a 1-directional DME within the same singular plane GW.
There are also exact plane gravitational waves with a $\frac{1}{U^{2}}$ singularity near the origin. The simplest example is the one with the profile $K=\frac{k}{U^{2}}\rm{diag}(1,-1)$ whose symmetries worked in [28]. There is also the Lukash gravitational wave [29, 30], which is endowed with an extra symmetry but suffers from the same singularity [^6] Same singularity appears in nuclear physics, see [31] and references therein..
Now, we show that SUSY QM allows us to seek DME within such singular waves. The approximated profile can be constructed out of the superpotential
@@
which is singular at the origin. Partner potentials, in this case, are
$V_{1}$ $=A^{2}+A(A+1)\operatorname{csch}^{2}U;$ (5.2a)
@@
This singular potential is shape invariant with $A\to A-1$; thus integrable. Note that $W(U)$ (5.1) is parity odd; e.g., $W(\infty)=-W(-\infty)=1$. In order to obtain a non-vanishing DME, we select $A=n$, where $n$ indexes the bound states. Then, the associated geodesic equation becomes
@@
$X_{1}$ $=\coth U;\qquad\qquad\ \ \Delta X_{1}=2;$ (5.4a)
$X_{2}$ $=3\coth^{2}U-1;\qquad\Delta X_{2}=0.$ (5.4b)
These solutions blow up at $U=0$, as expected. Using the parity argument given in (3.11), we deduce that only geodesics indexed with odd $n$ have non-vanishing DME. For instance, we obtain $\Delta X_{3}=12$ without even solving for $X_{3}(U)$.
If we were to build a plane wave including the partner potential $V_{2}$, it would be pretty much similar to the PT plane wave described in Section (4.2). On the other hand, if we build an exact plane GW with $K$ being
$K_{sing}=\begin{pmatrix}n(n+1)\operatorname{csch}^{2}U&0\\ 0&-n(n+1)\operatorname{csch}^{2}U\,\end{pmatrix}$ (5.5)
where we obtain DME only in the $X$-sector.
Based on our model, we conclude that DME can also be found in GWs endowed with $1/U^{2}$ singularity. Of course, memory effects in singular profiles should be considered with more care. Our arguments here can be taken as a necessary first step towards a complete analytic solution.

### v3

@@ 3.3 DME of plane GWs
In this section, we build the related geodesics using SUSY methods and investigate their DME. Our calculations are made for $d=1$ case. However, they are valid for $d=2$ plane and GWs of Section 4 as their profiles $K_{ij}(U)$ are diagonal.
Let us begin with $n=1$. Using the ladder operator $L^{\dagger}$ (3.7), we get
@@
Recall that $A$ is the shape invariance parameter, thus it cannot be equal to 0. The negative values of $A$ yield diverging geodesics with diverging velocities at infinity. Thus, $X_{1}(U)$ satisfies the DME condition (2.5) provided that $A\geq 1$  [^4] It is a half wave in the language of [14, 15, 16].. One can also see that as $U\to\pm\infty$, $X_{1}(U)\neq 0$ only if $A=n=1$.
Subsequent SUSY Darboux transformations (3.7) generate other geodesics with higher half-wave numbers $n$.
With this unified approach, we will show below that the PT and the Scarf profiles in [14, 15, 16] are actually two different faces of the same theory, viewed from the perspective of SUSY QM. Both profiles are rooted in (3.9), hence derivable from (3.10).
For this, an important point is the value of the free parameter $A\geq n$. If we set $A=n$, the formula for geodesics (3.8) simplifies in terms of $W(U)$ as
@@
Substituting (3.16) into (3.17), we obtain
$\Delta X_{n}=W(\infty;n)X_{n-1}(\infty;n-1)-W(-\infty;n)X_{n-1}(-\infty;n-1),$ (3.18)
as our geodesics satisfy the Neumann boundary condition (2.5). Therefore, SUSY QM yields a compact and useful formula (3.18) for the DME of plane waves in terms of the asymptotic values of superpotential and the preceding geodesics [^5] Our formula (3.18) might be somewhat reminiscent of Christodoulou’s nonlinear memory [31].. Via (3.17) we observe that a non-vanishing DME requires parity odd geodesics and our second formula (3.18) relates this condition to the parity of the superpotential. But the individual parts of $W(U)$ (3.9) have different parity properties. Thus, we investigate each part separately and show that these magical numbers come out of our unified picture in a beautiful manner.
3.3.1 The Pöschl-Teller profile
To recover the PT-type profile (2.6), we need to set
@@ 3.3.1 The Pöschl-Teller profile
$A-n$ $=0,$ (3.19b)
in (3.13). While the first condition above eliminates the parity-odd Scarf term, the second one selects the highest excited state by quantizing the parameter $A$. An integer $A$ leads to the so-called reflectionless PT problem. In that case, the geodesic equation turns out to be
@@
cf. (2.2a). As we are dealing with a shape invariant profile, the bounded geodesics can be found via (3.7), where we set $A=m$ first,
@@
where $X_{0}(U,m)=\operatorname{sech}^{m}U$ and $m=1,2,3,\cdots$.
@@
For example, if we take $n=1$, we obtain
@@
$X_{1}=\tanh U,$ (3.23)
which is consistent with [14, 15]. The related DME can be computed easily from (3.18) as
@@
Similarly, $n=2$ geodesic is obtained using (3.21) as
@@
For $n=3$, the solution of the geodesic equation (3.20) is
$X_{3}(U;m)=(2m-5)(2m-3)\Big{(}2(m-2)\operatorname{sech}^{m-3}U+(1-2m)\operatorname{sech}^{m-1}U\Big{)}\tanh U.$ (3.27)
@@
With SUSY, one can go on and build geodesics for any $n$ and we observe that $\Delta X\neq 0$ only if $m=n$ and $n$ is an odd integer. We conclude that the geodesics endowed with DME are nothing but the highest excited states of the analogous quantum mechanical problem with Neumann boundary condition.
Longitudinal motion and memory:
To compute the $V$-component of the geodesics for the first two values of $n$, we substitute our solutions $X(U)$ i.e., (3.23) and (3.26) into (2.3), respectively. We observe that part of the longitudinal motion organizes itself as another particular solution of the PT problem. For instance,
@@
$V_{2}(U)$ $=$ $-3\tanh U\operatorname{sech}^{2}U(2-3\operatorname{sech}^{2}U)+c_{3}U+c_{4}=-\frac{1}{35}X_{3}(U;m=5)+c_{3}U+c_{4},\qquad\quad$ (3.30)
where $c_{1},..,c_{4}$ are constants. $X_{1}(U;m=3)$ and $X_{3}(U;m=5)$ can be found from (3.22) and (3.27), respectively. These motions exhibit VME when the initial constants $c_{1},c_{3}$ do not vanish.
3.3.2 The Scarf profile
In order to obtain the geodesic equation for the Scarf profile (2.7), we need to satisfy another set of equations
@@ 3.3.2 The Scarf profile
in (3.13). While the first equation quantizes the parameter $A$, (3.31b) solves for $B$ non-trivially as
@@
where $n=0,1,2,\cdots$. Pay attention at that $B$ has double roots as in [16]. This multiplicity will allow us to extend our arguments to exact plane gravitational waves with 2-transverse directions [^6] That extension was first done in [16]. Here, we elaborate on their results using SUSY methods. in Section 4.
When $d=1$ (3.13) reduces to a geodesic equation (2.2)
@@
augmented with the boundary condition (2.5). Thus, supersymmetry leads to a quick and elegant derivation of the Scarf geodesic equation which mimics the derived Gaussian flyby profile. In [16] the above equation was obtained through the hypergeometric equation and solved with the Nikiforov-Uvarov method. SUSY provides a more economical way to find these solutions.
@@
The seed geodesic can be found via (3.4) as
$X_{0}(U)=\operatorname{sech}^{m}U\ e^{-B\tan^{-1}(\sinh U)},$ (3.34)
where $A=m$ and $B=\pm\sqrt{m(m+1)}$. As before, $m$ is a positive integer. When $n=1$, a family of geodesics labeled by $m$ can be obtained from (3.7) as
@@
(3.35) yields a non-vanishing DME only when $m=n=1$. In this case, it reduces to
@@
Let us emphasize that (3.36) satisfies (3.33) with $n=1$. The related DME can be computed either directly from the geodesic solution above or from (3.18)
$\Delta X_{1}=2\cosh\left(\frac{B\pi}{2}\right)=2\cosh\left(\frac{\pi}{\sqrt{2}}\right),\quad B=\pm\sqrt{2},$ (3.37)
@@
Applying a second Darboux transformation, we generate $X_{2}(U;m)$. In order to satisfy (3.33), we set $m=2$ and get
@@
By virtue of the supersymmetric transformations (3.16), one may obtain $n\geq 3$ geodesics with non-zero DME in a straightforward manner.
We would like to emphasize the non-trivial quantization of $B$ (3.32) in the Scarf profile. This sign ambiguity makes it totally different from the PT case where the sole parameter $A$ takes only positive integer values. Another difference is that the Scarf profile may yield a non-vanishing DME even for $m\neq n$. The same results can be obtained via the second potential $V_{2}$ (3.10b) by shifting $A\to A+1$.
Longitudinal motion and memory:
In order to illustrate the longitudinal memory for the Scarf profile, we construct $V_{1}(U)$ (2.3) as
@@
Since $\frac{dV_{1}}{dU}|_{U\to\pm\infty}=c_{1},$ the coordinate $V$ exhibits a VME unless $c_{1}$ vanishes. This result can be extended to any $V_{n}(U)$ with $n\geq 2$.
@@
Those $d=1$ cases studied above can be extended to plane waves with 2-transverse directions. For this, we choose a diagonal profile $K_{ij}(U)$ in (2.1) and set $X^{i}=\{X,Y\}$.
4.1 Exact plane gravitational wave with the Scarf profile
We begin with the Scarf case. As in [16], the double root property of $B$ (3.32) allows us to construct an exact plane GW (vacuum solution of Einstein’s equations) mimicking flyby. The associated profile is diagonal and traceless
$K_{scarf}=\begin{pmatrix}(2n+1)|B|\operatorname{sech}U\tanh U&0\\ 0&-(2n+1)|B|\operatorname{sech}U\tanh U\,\end{pmatrix},$ (4.1)
as $B=\pm|B|$ and $|B|=\sqrt{n(n+1)}$ with $n$ being a positive integer. As we discussed before, the $n=1$ and $n=2$ geodesics with their DMEs can be simply extracted from Section 3.3.2 as
@@ 4.1 Exact plane gravitational wave with the Scarf profile
$Y_{1}(U)$ $=\left[\tanh U-2\sqrt{2}\operatorname{sech}U\right]e^{\sqrt{2}\tan^{-1}(\sinh U)},\qquad\quad\Delta Y_{1}=2\cosh\left(\frac{\pi}{\sqrt{2}}\right),$ (4.2b)
and
@@
$Y_{2}$ $=\left[2+21\operatorname{sech}^{2}U-8\sqrt{6}\tanh U\operatorname{sech}U\right]e^{\sqrt{6}\tan^{-1}(\sinh U)},\qquad\Delta Y_{2}=4\sinh\Big{(}\frac{\sqrt{6}\pi}{2}\Big{)}.$ (4.3b)
These results are consistent with [16].
It is also possible to find lightcone geodesic $V_{n}(U)$ by combining solutions of those transverse geodesics above. As explained in Section 3.3.2, the longitudinal memory will be a VME.
We may construct further plane wave solutions where $m>n$, see (3.35). In that case, the trace of the profile is non-zero, i.e., $\mathrm{Tr}{K}=2(m-n)^{2}$ (3.13). One can get a non-vanishing DME when, for instance, $n=2$ and $m=3$.
We emphasize that the construction of the exact plane GW (4.1) with $\mathrm{Tr}(K)=0$ (4.1) depends heavily on the non-trivial quantization of the parameter $B$ with sign ambiguity. Based on this, one can combine the terms with different signs. This is no longer true for the PT plane waves as it will be outlined below.
@@ 4.2 Pöschl-Teller plane waves
In [15], it was argued that when an exact plane GW is built out of a PT profile, DME can be obtained only in one direction. The other direction should be put to zero via appropriate initial conditions. SUSY provides an argument for this: Unlike as in the Scarf case, the parameter $A$ of the PT profile has a single root (3.19b) which can take only positive integer values. The shape invariance condition allows us to build a family of profiles labeled by $n$ but always with the same sign.
To understand this better, we construct a 4-dimensional plane wave out of the supersymmetric partners (3.10). The result will be a null-fluid spacetime, $\mathrm{Tr}(K)\neq 0$, with the following diagonal profile
$K_{PT}=\begin{pmatrix}(m-n)^{2}-m(m+1)\operatorname{sech}^{2}U&0\\ 0&(m-n)^{2}-m(m-1)\operatorname{sech}^{2}U\,\end{pmatrix},$ (4.4)
where parameters $m$ and $n$ are integers. From Section 3.3.1, we recall that DME occurs only when $m=n$. That simplifies (4.4) and puts it into its final form. Then, the resulting geodesic equations (2.2a) are
@@
$\left(\frac{d^{2}}{dU^{2}}+m(m-1)\operatorname{sech}^{2}U\right)Y$ $=0.$ (4.5b)
Due to shape invariance, the solutions $X(U)$ and $Y(U)$ are interrelated: they match with the exception for the ground state of $X$. If one sets $m=1$, the solution for the first one is $X=\tanh U$ (3.23). The second equation, on the other hand, becomes that of a free particle and confirms the claim made in [15]. Recall that choosing $m$ as an integer leads to a reflectionless PT potential : the reflection coefficient vanishes just like for a free particle. The shape invariance integrability condition (3.11) puts a limitation on the parameter $A$ so that it takes only positive values. Therefore, PT plane wave profiles (4.4) and (4.5) always share the same sign and it is impossible to build an exact plane GW in that case.
Other choices of $m>1$ yield DME again in one direction. For instance, when $m=2$, $X=3\tanh^{2}U-1$ and $Y=\tanh U$. Therefore, using superpartners in the profile yields a tower of geodesics exhibiting DME in alternating directions.
@@ 5 Modeling singular profiles with SUSY
Our final goal is to show that SUSY provides further insight and predicts DME for other GWs.
In [32, 33], memory effects of the profile $K(U)=U^{-4}$ were discussed. Later, the authors in [12] obtained a DME within the same singular plane GW.
There exist also exact plane GWs with a $U^{-2}$ singularity near the origin. The simplest example is the one with profile
$K(U)=\frac{k}{U^{2}}\rm{diag}(1,-1),$ (5.1)
whose symmetries were studied in [34]. The Lukash GW [35, 36], which is endowed with an extra symmetry suffers from the same singularity [^7] That $\frac{1}{U^{2}}$ singularity appears in nuclear physics too, see [37] and references therein..
Now, we show that SUSY QM allows us to search for DME also within such singular plane waves. The related profile can be constructed using the following superpotential
@@
which is singular at $U=0$. In this case the partner potentials become
$V_{1}$ $=A^{2}+A(A+1)\operatorname{csch}^{2}U,$ (5.3a)
@@
They are shape invariant with $A\to A-1$, thus the system is integrable. Note that the superpotential (5.2) is parity odd; e.g., $W(\infty)=-W(-\infty)=1$. In order to obtain a non-vanishing DME, we select $A=n$, where $n$ labels the geodesics. Then, the associated geodesic equation becomes
@@
$X_{1}$ $=\coth U,\qquad\qquad\ \ \Delta X_{1}=2,$ (5.5a)
$X_{2}$ $=3\coth^{2}U-1,\qquad\Delta X_{2}=0.$ (5.5b)
These solutions blow up at $U=0$, as expected. Using the parity argument given in (3.17) and (3.18), we deduce that only geodesics labeled with odd $n$ have a non-vanishing DME. For instance, we obtain $\Delta X_{3}=12$ without even solving for $X_{3}(U)$.
If we were to build a plane wave including the partner potential $V_{2}$, it would be pretty much similar to the PT plane wave described in Section (4.2). On the other hand, if we build an exact plane GW with
$K_{sing}=\begin{pmatrix}n(n+1)\operatorname{csch}^{2}U&0\\ 0&-n(n+1)\operatorname{csch}^{2}U\,\end{pmatrix},$ (5.6)
we obtain DME only in the $X$-sector with the same argument as in Section 4.2.
Based on our model (5.3), we conclude that DME can also be found in GWs with $U^{-2}$ singularity. Of course, the memory effects in singular profiles should be handled with more care. Our arguments here can be taken as a necessary first step towards a complete solution which will be reported elsewhere.